# Token Reduction Should Go Beyond Efficiency in Generative Models – From Vision, Language to Multimodality

### v2

@@ Abstract
In Transformer architectures, tokens—discrete units derived from raw data—are formed by segmenting inputs into fixed-length chunks. Each token is then mapped to an embedding, enabling parallel attention computations while preserving the input’s essential information. Due to the quadratic computational complexity of transformer self-attention mechanisms, token reduction has primarily been used as an efficiency strategy. This is especially true in single vision and language domains, where it helps balance computational costs, memory usage, and inference latency. Despite these advances, this paper argues that token reduction should transcend its traditional efficiency-oriented role in the era of large generative models. Instead, we position it as a fundamental principle in generative modeling, critically influencing both model architecture and broader applications. Specifically, we contend that across vision, language, and multimodal systems, token reduction can: (i) facilitate deeper multimodal integration and alignment, (ii) mitigate "overthinking" and hallucinations, (iii) maintain coherence over long inputs, and (iv) enhance training stability, etc. We reframe token reduction as more than an efficiency measure. By doing so, we outline promising future directions, including algorithm design, reinforcement learning-guided token reduction, token optimization for in-context learning, and broader ML and scientific domains. We highlight its potential to drive new model architectures and learning strategies that improve robustness, increase interpretability, and better align with the objectives of generative modeling.  [^1] We collected a list of token reduction papers at: Awesome-Collection-Token-Reduction.
@@ 1 Introduction
Transformer-based generative models brown2020language ; devlin2019bert ; NEURIPS2021_854d9fca ; vaswani2017attention have emerged as dominant deep learning architectures across vision, language, and multimodal tasks, due to their ability to process long sequences of tokens, which are the fundamental representational units derived from raw data such as subwords in language or image patches in vision. As these models are applied to increasingly complex real-world tasks, the input sequence lengths of both the models and their training datasets continue to grow. However, the quadratic computational complexity of the attention mechanism results in high memory usage and slow inference, which hinders the practical deployment of generative models at scale. Token reduction addresses this challenge by reducing the number of tokens processed during inference. By pruning or merging tokens, token reduction decreases computational cost and accelerates runtime, offering a practical solution for improving the efficiency of generative models.
Token reduction has been widely adopted in computer vision, language processing, and multimodal tasks. In vision transformers, it has primarily been used to reduce computational cost by removing visually redundant tokens bolya2022token ; kong2022spvit ; liang2022not ; rao2021dynamicvit . While effective in decreasing token-level computations and accelerating inference, this approach can discard subtle but important visual features, especially in dense prediction tasks such as segmentation and object detection, leading to significant performance degradation. In language models, token reduction has commonly been implemented through early-exit mechanisms and token-skipping strategies lin2025boosting ; wu2024accelerating , which reduce the number of intermediate tokens processed and thus lower computational overhead. Similarly, multimodal large language models (MLLMs) apply visual token pruning primarily during the prefill stage chen2024image , where adaptive attention patterns are learned in the early layers to prune tokens in later stages. Despite progress in existing work, token reduction is still predominantly viewed as a means of improving computational efficiency, primarily by reducing the number of tokens to minimize associated computations and accelerate inference. Such an efficiency-only mindset has critical limitations. Naive pruning methods may discard informative tokens, thereby degrading model understanding and performance liang2022not ; zhan2024exploring ; zhan-etal-2024-rethinking-token . Furthermore, token reduction is commonly treated as a post hoc optimization, rather than being integrated into the core design and training of the model chen2024image .
@@
Modern generative tasks present numerous challenges that highlight the need for thoughtful token selection: (i) Ultra-long contexts in language modeling require selective retention of relevant segments to preserve coherence. (ii) LLMs frequently exhibit overthinking, repeatedly attending to low-value tokens and producing redundant or contradictory outputs. (iii) Multimodal generation tasks often face issues of visual redundancy, where background tokens overshadow salient visual features critical for accurate understanding. (iv) Noisy or irrelevant tokens introduced during training slow down convergence and harm model stability. By learning to intelligently select, merge, or compress tokens based on their contribution to generation objectives-rather than solely on raw redundancy-models can simultaneously reduce computational load, improve robustness, and enhance interpretability and alignment. This paper makes the following three key contributions:
  • We provide a comprehensive analysis of the limitations inherent in current efficiency-centric token pruning methods across vision, language, and multimodal generative models, highlighting critical shortcomings in these widely adopted strategies.
@@
  • We propose a roadmap for future research on token reduction, including directions for method design, reinforcement learning-guided token selection, adaptive in-context compression, and hardware-algorithm co-design, etc. These directions aim to support the development of next-generation generative architectures that are both robust and efficient.
@@ 2.1 Token Reduction in Vision Models
Image Classification. Classification serves as a fundamental task for vision models and token reduction techniques have been widely applied in it due to its simplicity and versatility. It has been widely explored from various aspects bolya2022token ; liang2022not ; rao2021dynamicvit ; wu2023ppt ; zeng2024m2m ; kong2023peeling ; zeng2025token . Specifically, DynamicViT rao2021dynamicvit devises a lightweight module to predict the importance score of each token, thereby pruning unimportant tokens. SPViT kong2022spvit introduces a soft pruning technique, which integrates the less informative tokens generated by the selector module into a package token that will participate in subsequent calculations rather than being completely discarded. EViT liang2022not identifies attentive tokens from the attention map, enabling token pruning without additional parameters. ToMe bolya2022token merges tokens with similarity based on bipartite matching to maintain information utility. PPT wu2023ppt analyzes the statistic data between layers and adaptively employs token pruning and merging within layers to achieve higher acceleration performance.
@@
Year2020 - 2021Power-bert goyal2020power , DynamicViT rao2021dynamicvit ,TokenLearner ryoo2021tokenlearner , SpAtten wang2021spatten 2022LTP kim2022learned , SPViT kong2022spvit , EViT liang2022not ,PatchSlim tang2022patch , STTS wang2022efficient , PPT ma2022ppt 2023DCP anagnostidis2023dynamic , ToMe bolya2022token , ToMeSD bolya2023token , PuMer cao2023pumer ,HeatViT dong2023heatvit , Llmlingua jiang2023llmlingua ,DTP nawrot2022efficient , TPS wei2023joint 2024Honeybee cha2024honeybee , FastV chen2024image , TokenPacker li2024tokenpacker ,BRAVE kar2024brave , LLaMA-VID li2024llama , DyDiT zhao2024dynamic - 2025.07DivPrune alvar2025divprune , SeTok wu2024towards , TopV yang2025topv , SparseVLM zhang2024sparsevlm Voco-llama ye2024voco , S2-MAD zeng2025s , ToCa zou2025accelerating ,TALE han2024token Vision & Language TransformerLLM & Multimodal LLMEfficiency
$\bm{\checkmark}$
$\bm{\Rightarrow}$  Future ?
@@
Figure 1: Timeline of notable method developments for token reduction methods with modality shifts (Vision & Language $\to$ Multimodal LLMs). All these strategies aim to speed up inference with negligible performance drops. Conversely, we ask: What is the next token reduction paradigm in generative model design that goes beyond test-time accelerations?
@@
Video Compression. Unlike token reduction in image classification, video compression focuses more on the temporal redundancy within videos, and algorithms are developed to reduce the number of tokens with less computational overhead. Various token reduction methods have been investigated for different tasks, including video understanding ryoo2021tokenlearner ; sun2025llavascissortokencompressionsemantic ; wang2022efficient , video editing li2024vidtome , video-text retrieval liu2022ts2 ; shen2024tempme , video action detection chen2023efficient , and so on. Specifically, STTS wang2022efficient introduces a lightweight framework that dynamically selects the most informative spatial-temporal tokens in video transformers. Tokenlearner ryoo2021tokenlearner proposes an adaptive tokenization module that learns a handful of informative spatial-temporal tokens, significantly reducing computational costs. EVAD chen2023efficient selectively drops irrelevant spatial-temporal tokens in non-keyframes while preserving keyframe and motion-relevant tokens, and then refines actor features using a context-aware decoder to maintain accuracy with reduced computations.
@@
Generative Tasks. Token reduction in generative tasks ju2024turbo aims to accelerate generative models through the efficient utilization of tokens. It can be applied to both diffusion models bolya2023token ; wang2024attention and diffusion transformers zou2025accelerating . Specifically, ToMeSD bolya2023token exploits natural redundancy in generated images by merging redundant tokens, successfully extending token merging to stable diffusion with simple unmerging. DyDiT zhao2024dynamic reduces redundancy with a Timestep-wise Dynamic Width approach to adopt model width conditioned on the generation timesteps, and a Spatial-wise Dynamic Token strategy to avoid redundant computations at unnecessary spatial locations.
@@ 2.2 Token Reduction in Language Models
Token reduction strategies in language modeling have evolved from early optimizations for BERT huang2022pyramid ; kim2020length ; kim2022learned ; kim2023leap ; ye2021tr to techniques specifically designed for LLMs. PoWER-BERT goyal2020power introduces progressive word-vector elimination by removing redundant token representations based on self-attention dynamics, improving inference efficiency. Learned token pruning kim2022learned extends this approach by learning attention-based thresholds to adaptively prune uninformative tokens, thereby reducing computational costs while preserving model performance. In LLMs, token reduction must account for the constraints of autoregressive decoding across diverse downstream tasks. Dynamic pooling methods anagnostidis2023dynamic ; tao2025saliency adjust token representations on the fly during inference to reduce redundancy. Prompt compression techniques fu2024lazyllm ; ge2023context ; jiang2023llmlingua aim to reduce computational overhead by compressing the input prompt before generation. Selective decoding approaches fu2024lazyllm ; wingate2022prompt reduce per-step inference costs by computing key-value pairs only for tokens critical to predicting the next token. In multi-agent systems, S2-MAD zeng2025s proposes a sparsification mechanism that limits unnecessary token exchanges between agents, reducing communication costs and improving the efficiency of collaborative reasoning.
@@ 2.3 Token Reduction in Multimodal LLMs
Recent work has explored visual token pruning by addressing attention inefficiencies of deep transformer layers in MLLMs alvar2025divprune ; arif2025hired ; lin2025boosting ; shang2024llava . Specifically, FastV chen2024image shows that deeper vision-language layers expend significant computations on redundant image tokens. To address this, a lightweight module is adopted to adaptively prune these tokens, reducing inference overheads in subsequent stages. A complementary approach modifies vision feature extractors or projectors to output a smaller set of highly informative image tokens, effectively distilling the input into a compressed representation cai2024matryoshka ; cha2024honeybee ; kar2024brave ; li2024tokenpacker ; li2024mini ; ye2024voco . However, efficiency gains from these prefill stages often fade during the decoding phase, where per-token computations dominate. To overcome this, recent methods jointly optimize token reduction during both prefill and decoding stages, ensuring sustained speedups throughout inference huang2024dynamic ; song2024less . In Fig. 1, we present a timeline of notable developments in token reduction methods, illustrating the shift from early applications in ViT and BERT-based models to more recent advances in LLMs and MLLMs.
@@ 3 Problem Formulation
In modern generative models Qwen2.5VL ; 2024llama3 ; liu2023llava ; openai2024gpt4 ; Peebles2022DiT , a token denotes one fundamental unit of input or representation, typically encoded as a vector. For example, a token might correspond to a subword in language, a patch in an image, or an embedding of a time step in audio. We denote a sequence of $N$ input tokens as $X=[x_{1},\dots,x_{N}]\in\mathbb{R}^{d}$. Token reduction refers to any operation that compresses the token sequence to $M$ tokens (with $M<N$) by removing or consolidating tokens while aiming to preserve the original information.
Broadly, token reduction methods fall into the following categories: 1) Token pruning methods bolya2022token that remove entire unimportant tokens, simply dropping them from the sequence; 2) Token merging methods bolya2022token ; bolya2023token which fuse information from multiple tokens into fewer tokens, effectively compressing the sequence by merging similar or related tokens; 3) Hybrid strategies cao2023pumer ; kim2024token ; wu2023ppt that combine pruning and merging within a unified framework; 4) Token distillation approaches cai2024matryoshka ; mu2023learning which integrate rich information across longer input sequences or multiple modalities into fewer condensed tokens, enabling efficient cross-modal interactions and long-context reasoning in LLMs and MLLMs.
A core challenge in token reduction is the determination of tokens to be pruned or merged. There are various importance criteria and scoring mechanisms to rank token significance, including attention-based heuristics liang2022not , gradient or loss-based criteria huang2024dynamic , clustering haurum2024agglomerative , and learned predictors rao2021dynamicvit .
From a purely efficiency-oriented perspective, token reduction delivers substantial computational efficiency gains by reducing the quadratic computation cost from $\mathcal{O}(N^{2})$ to $\mathcal{O}(M^{2})$ in attention mechanisms. By eliminating redundant tokens and processing fewer computations during inference, it effectively accelerates the inference speed and improves the model throughput, which is crucial for latency-sensitive tasks or real-time applications. Furthermore, it also shrinks the memory footprint for activations and gradients (e.g., key/value caches), alleviating memory usage for both inference and training, which is particularly beneficial for wide deployments on resource-limited platforms.
@@
4 Challenges and Positions
In this section, we establish token reduction as a foundational mechanism for addressing critical challenges in modern generative systems. We claim five core challenges across modalities: visual representation sparsity, semantic misalignment, reasoning redundancy, training instability, and long-context overload. We demonstrate how principled token reduction strategies intrinsically address these issues through dynamic token-semantic co-optimization. We position token reduction not only as an efficiency tool, but as an essential paradigm for enhancing semantic coherence and enabling sustainable scaling of generative systems.

### v3

@@ Abstract
In Transformer architectures, tokens—discrete units derived from raw data—are formed by segmenting inputs into fixed-length chunks. Each token is then mapped to an embedding, enabling parallel attention computations while preserving the input’s essential information. Due to the quadratic computational complexity of transformer self-attention mechanisms, token reduction has primarily been used as an efficiency strategy. This is especially true in single vision and language domains, where it helps balance computational costs, memory usage, and inference latency. Despite these advances, this paper argues that token reduction should transcend its traditional efficiency-oriented role in the era of large generative models. In this paper, we characterize this mechanism as a fundamental principle in generative modeling, critically influencing both model architecture and broader applications. We analyze how token reduction addresses critical challenges in current systems across vision, language, and multimodal, demonstrating its ability to: (i) facilitate deeper multimodal integration and alignment, (ii) mitigate "overthinking" and hallucinations, (iii) maintain coherence over long inputs, and (iv) enhance training stability, etc. We reframe token reduction as more than an efficiency measure. By doing so, we outline promising future directions, including algorithm design, reinforcement learning-guided token reduction, token optimization for in-context learning, agentic framework design, and broader ML and scientific domains.  [^1] We collected a list of token reduction papers at: Awesome-Collection-Token-Reduction.
@@ 1 Introduction
Transformer-based generative models Brown et al. (2020); Devlin et al. (2019); Han et al. (2021); Vaswani et al. (2017) have emerged as dominant deep learning architectures across vision, language, and multimodal tasks, due to their ability to process long sequences of tokens, which are the fundamental representational units derived from raw data such as subwords in language or image patches in vision. As these models are applied to increasingly complex real-world tasks, the input sequence lengths of both the models and their training datasets continue to grow. However, the quadratic computational complexity of the attention mechanism results in high memory usage and slow inference, which hinders the practical deployment of generative models at scale. Token reduction addresses this challenge by reducing the number of tokens processed during inference. By pruning or merging tokens, token reduction Guo et al. (2025); Han et al. (2026); Huang et al. (2024a); Hyun et al. (2025); Jiang et al. (2025b); Kim et al. (2025); Li et al. (2025a); Liu et al. (2025b); Wu et al. (2025e); Zhang et al. (2025d) reduces computational cost and accelerates runtime, providing a practical solution for enhancing generative efficiency Kim et al. (2025); Liu et al. (2025c); Tang et al. (2025); Wu et al. (2025a); Xing et al. (2025); Yao et al. (2025); Zhang et al. (2025a); Zhang and Fu (2025); Zhang et al. (2025e); Zhuang et al. (2025).
Token reduction has been widely adopted in computer vision, language processing, and multimodal tasks. In vision transformers, it has primarily been used to reduce computational cost by removing visually redundant tokens Bergner et al. (2025); Bolya et al. (2023); Fang et al. (2025a); Hong and Liu (2025); Kong et al. (2022); Lei et al. (2025); Liang et al. (2022b); Rao et al. (2021). In language models, token reduction has commonly been implemented through early-exit mechanisms and token-skipping strategies Lin et al. (2025); Wu et al. (2024), which reduce the number of intermediate tokens processed and thus lower computational overhead. Similarly, multimodal large language models (MLLMs) apply visual token pruning primarily during the prefill stage Chen et al. (2024), where adaptive attention patterns are learned in the early layers to prune tokens in later stages. Despite progress, token reduction is predominantly viewed as a post-hoc efficiency optimization Liu et al. (2025d); Shao et al. (2025), primarily by reducing the number of tokens to minimize associated computations and accelerate inference. Such an efficiency-only mindset has limitations. Naive pruning methods may discard informative tokens, thereby degrading model understanding and performance Liang et al. (2022b); Zhan et al. (2024a, b). Furthermore, token reduction is commonly treated as a post hoc optimization, rather than being integrated into the core design and training of the model Chen et al. (2024).
@@
Modern generative tasks present numerous challenges that highlight the need for thoughtful token selection: (i) Ultra-long contexts in language modeling require selective retention of relevant segments to preserve coherence. (ii) LLMs frequently exhibit overthinking, repeatedly attending to low-value tokens and producing redundant or contradictory outputs. (iii) Multimodal generation tasks often face issues of visual redundancy, where background tokens overshadow salient visual features critical for accurate understanding. (iv) Noisy or irrelevant tokens introduced during training slow down convergence and harm model stability. By learning to intelligently select, merge, or compress tokens based on their contribution to generation objectives, rather than solely on raw redundancy, models can simultaneously reduce computational load, improve robustness, and enhance interpretability and alignment. This paper makes the following three key contributions:
  • We categorize existing token reduction methods by their functional objective, identifying a transition from efficiency-centric optimizations to task-aware enhancements in vision, language, and multimodal domains.
@@
  • We outline a roadmap for future research on token reduction, including directions for method design, reinforcement learning-guided token selection, adaptive in-context compression, and hardware-algorithm co-design, etc. These directions aim to support the development of next-generation generative architectures that are both robust and efficient.
@@ 2.1 Token Reduction in Vision Models
Image Classification. Classification serves as a fundamental task for vision models and token reduction techniques have been widely applied in it due to its simplicity and versatility. It has been widely explored from various aspects Bolya et al. (2023); Liang et al. (2022b); Rao et al. (2021); Wu et al. (2023); Zeng and Yu (2024); Kong et al. (2023); Zeng et al. (2025a). Specifically, DynamicViT Rao et al. (2021) devises a lightweight module to predict the importance score of each token, thereby pruning unimportant tokens. SPViT Kong et al. (2022) introduces a soft pruning technique, which integrates the less informative tokens generated by the selector module into a package token that will participate in subsequent calculations rather than being completely discarded. EViT Liang et al. (2022b) identifies attentive tokens from the attention map, enabling token pruning without additional parameters. ToMe Bolya et al. (2023) merges tokens with similarity based on bipartite matching to maintain information utility. PPT Wu et al. (2023) analyzes the statistic data between layers and adaptively employs token pruning and merging within layers to achieve higher acceleration performance.
@@
Year2020 - 2021Power-bert Goyal et al. (2020), DynamicViT Rao et al. (2021),TokenLearner Ryoo et al. (2021), SpAtten Wang et al. (2021)2022LTP Kim et al. (2022), SPViT Kong et al. (2022), EViT Liang et al. (2022b),PatchSlim Tang et al. (2022), STTS Wang et al. (2022), PPT Ma et al. (2022)2023DCP Anagnostidis et al. (2023), ToMe Bolya et al. (2023), ToMeSD Bolya and Hoffman (2023), PuMer Cao et al. (2023),HeatViT Dong et al. (2023), Llmlingua Jiang et al. (2023),DTP Nawrot et al. (2023), TPS Wei et al. (2023)2024Honeybee Cha et al. (2024), FastV Chen et al. (2024), LLaMA-VID Li et al. (2024c), TokenPacker Li et al. (2024a), BRAVE Kar et al. (2024), DyDiT Zhao et al. (2025), TALE Han et al. (2025)2025DivPrune Alvar et al. (2025), SeTok Wu et al. (2025b), SparseVLM Zhang et al. (2025g) TopV Yang et al. (2025a), Voco-llama Ye et al. (2025), FastDriveVLA Cao et al. (2025),EfficientVLA Yang et al. (2025c), S2-MAD Zeng et al. (2025b), LatentMAS Zou et al. (2025b)Vision & Language TransformerLLM & Multimodal LLM & AgentsEfficiency
[…]
@@
Figure 1: Timeline of notable method developments for token reduction methods with modality shifts (Vision & Language $\to$ Multimodal LLMs & Agents). All these strategies aim to speed up inference with negligible performance drops. Conversely, we ask: What is the next token reduction paradigm in generative model design that goes beyond test-time accelerations?
@@
Video Compression. Unlike token reduction in image classification, video compression focuses more on the temporal redundancy within videos, and algorithms are developed to reduce the number of tokens with less computational overhead. Various token reduction methods have been investigated for different tasks, including video understanding Ryoo et al. (2021); Sun et al. (2025); Wang et al. (2022), video editing Li et al. (2024b), video-text retrieval Liu et al. (2022); Shen et al. (2025a), video action detection Chen et al. (2023), and so on. Specifically, STTS Wang et al. (2022) introduces a lightweight framework that dynamically selects the most informative spatial-temporal tokens in video transformers. Tokenlearner Ryoo et al. (2021) proposes an adaptive tokenization module that learns a handful of informative spatial-temporal tokens, significantly reducing computational costs. EVAD Chen et al. (2023) selectively drops irrelevant spatial-temporal tokens in non-keyframes while preserving keyframe and motion-relevant tokens, and then refines actor features using a context-aware decoder to maintain accuracy with reduced computations.
@@
Generative Tasks. Token reduction in generative tasks Ju et al. (2024) aims to accelerate generative models through the efficient utilization of tokens. It can be applied to both diffusion models Bolya and Hoffman (2023); Lu et al. (2025); Wang et al. (2024a) and diffusion transformers Gao et al. (2023); Zou et al. (2025a). Specifically, ToMeSD Bolya and Hoffman (2023) exploits natural redundancy in generated images by merging redundant tokens, successfully extending token merging to stable diffusion with simple unmerging. DyDiT Zhao et al. (2025) reduces redundancy with a Timestep-wise Dynamic Width approach to adopt model width conditioned on the generation timesteps, and a Spatial-wise Dynamic Token strategy to avoid redundant computations at unnecessary spatial locations.
@@ 2.2 Token Reduction in Language Models
Token reduction strategies in language modeling have evolved from early optimizations for BERT Huang et al. (2022); Kim and Cho (2021); Kim et al. (2022, 2023); Ye et al. (2021) to techniques specifically designed for LLMs. PoWER-BERT Goyal et al. (2020) introduces progressive word-vector elimination by removing redundant token representations based on self-attention dynamics, improving inference efficiency. Learned token pruning Kim et al. (2022) extends this approach by learning attention-based thresholds to adaptively prune uninformative tokens, thereby reducing computational costs while preserving model performance. In LLMs, token reduction must account for the constraints of autoregressive decoding across diverse downstream tasks. Dynamic pooling methods Anagnostidis et al. (2023); Tao et al. (2025) adjust token representations on the fly during inference to reduce redundancy. Prompt compression techniques Fu et al. (2024); Ge et al. (2024); Jiang et al. (2023) aim to reduce computational overhead by compressing the input prompt before generation. Selective decoding approaches Fu et al. (2024); Wingate et al. (2022) reduce per-step inference costs by computing key-value pairs only for tokens critical to predicting the next token. In multi-agent systems, S2-MAD Zeng et al. (2025b) proposes a sparsification mechanism that limits unnecessary token exchanges between agents, reducing communication costs and improving the efficiency of collaborative reasoning.
@@ 2.3 Token Reduction in Multimodal LLMs
Recent work has explored visual token pruning by addressing attention inefficiencies of deep transformer layers in MLLMs Alvar et al. (2025); Arif et al. (2025); Lin et al. (2025); Shang et al. (2024). Specifically, FastV Chen et al. (2024) shows that deeper vision-language layers expend significant computations on redundant image tokens. To address this, a lightweight module is adopted to adaptively prune these tokens, reducing inference overheads in subsequent stages. A complementary approach modifies vision feature extractors or projectors to output a smaller set of highly informative image tokens, effectively distilling the input into a compressed representation Cai et al. (2025); Cha et al. (2024); Kar et al. (2024); Li et al. (2024a, d); Ye et al. (2025). However, efficiency gains from these prefill stages often fade during the decoding phase, where per-token computations dominate. To overcome this, recent methods jointly optimize token reduction during both prefill and decoding stages, ensuring sustained speedups throughout inference Huang et al. (2025); Song et al. (2025). Furthermore, the scope of token reduction has recently expanded to Vision-Language-Action (VLA) models, optimizing efficiency for real-time robotic manipulation and autonomous driving Cao et al. (2025); Jiang et al. (2025c); Pei et al. (2025); Yang et al. (2025c).
In Fig. 1, we present a timeline of notable developments in token reduction methods, illustrating the shift from early applications in ViT and BERT-based models to more recent advances in LLMs, MLLMs and Agent systems.
@@ 3 Problem Formulation
In modern generative models Bai et al. (2025); Grattafiori et al. (2024); Liu et al. (2023); OpenAI et al. (2024); Peebles and Xie (2023), a token denotes one fundamental unit of input or representation, typically encoded as a vector. For example, a token might correspond to a subword in language, a patch in an image, or an embedding of a time step in audio. We denote a sequence of $N$ input tokens as $X=[x_{1},\dots,x_{N}]\in\mathbb{R}^{d}$. Token reduction refers to any operation that compresses the token sequence to $M$ tokens (with $M<N$) by removing or consolidating tokens while aiming to preserve the original information.
Broadly, token reduction methods fall into four categories: 1) Token pruning methods Bolya et al. (2023) that remove entire unimportant tokens, simply dropping them from the sequence; 2) Token merging methods Bolya et al. (2023); Bolya and Hoffman (2023) which fuse information from multiple tokens into fewer tokens, effectively compressing the sequence by merging similar or related tokens; 3) Hybrid strategies Cao et al. (2023); Kim et al. (2024); Wu et al. (2023) that combine pruning and merging within a unified framework; 4) Token distillation approaches Cai et al. (2025); Mu et al. (2023) which integrate rich information across longer input sequences or multiple modalities into fewer condensed tokens, enabling efficient cross-modal interactions and long-context reasoning in LLMs and MLLMs.
A core challenge in token reduction is the determination of tokens to be pruned or merged. There are various importance criteria and scoring mechanisms to rank token significance, including attention-based heuristics Liang et al. (2022b), gradient or loss-based criteria Huang et al. (2025), clustering Haurum et al. (2024), and learned predictors Rao et al. (2021).
From a purely efficiency-oriented perspective, token reduction delivers substantial computational efficiency gains by reducing the quadratic computation cost from $\mathcal{O}(N^{2})$ to $\mathcal{O}(M^{2})$ in attention mechanisms. By eliminating redundant tokens and processing fewer computations during inference, it effectively accelerates the inference speed and improves the model throughput, which is crucial for latency-sensitive tasks or real-time applications. Furthermore, it reduces the memory footprint of activations and gradients (e.g., key/value caches), alleviating memory usage for both inference and training, which is particularly beneficial for wide-scale deployments on resource-limited platforms. We present more theoretical detail in the Appendix A
@@
4 Core Roles and Challenges
In this section, we discuss token reduction as a foundational mechanism for addressing critical challenges in modern generative systems. We categorize five core challenges across modalities: visual representation sparsity, semantic misalignment, reasoning redundancy, training instability, and long-context overload. We demonstrate how principled token reduction strategies intrinsically address these issues through dynamic token-semantic co-optimization. We position token reduction not only as an efficiency tool, but as an essential paradigm for enhancing semantic coherence and enabling sustainable scaling of generative systems.